# Solvability of the Korteweg-de Vries equation under meromorphic initial conditions by quadrature

### v1

@@ 1. Introduction
The outline of this paper is as follows. In Section 2 we collect preliminary results for proving Theorems 1.3 and 1.4. We give proofs of Theorems 1.3, 1.4 and 1.6 in Sections 3, 4 and 5, respectively. The result of Kovacic [13], which plays a crucial role in Section 5, is briefly reviewed in Appendix A.
@@ 2. Preliminary Results
The conclusion in Proposition 2.1(iv) holds under a weaker condition. For example, if $\lim_{k\to 0}b(k)b(-k)\to b_{0}$ $(\neq-1)$, then it holds.
@@
The system (1.5) is integrable in the sense of differential Galois theory unless $G$ is of type (vi). Proposition 2.3 plays a key role in the proof of Theorem 1.4 in Section 4.
@@
where $\tilde{u}(y)=u(1/y)$. Since $(k^{2}-\tilde{u}(y))/y^{4}$ has a pole of fourth order at $y=0$, we see by a standard result on higher-order scalar linear differential equations (see, e.g., Theorem 7 of [6]) that Eq. (1.5) has an irregular singularity at infinity $x=\infty$. To obtain a necessary condition for its integrabilty, we compute its formal monodromy, exponential torus and Stokes matrices [20, 22, 25] around $y=0$ for (4.1). See these references for necessary information in the following computations.
@@
$\frac{\mathrm{d}\eta}{\mathrm{d}y}+\eta^{2}+\frac{2}{y}\eta+\frac{k^{2}-\tilde{u}(y)}{y^{4}}=0$ (4.2)
@@
Substitution of $re^{i\theta}$ into (4.4) yields
$V(re^{i\theta})=\tilde{V}(re^{i\theta})e^{Q(re^{i\theta})},$
@@
$M=V(r)^{-1}V(re^{2\pi i})=\mathrm{id}$
by $V(re^{2\pi i})=V(r)M$. The exponential torus is given by
$\mathscr{T}=\mathrm{Gal}(\mathbb{C}(e^{\mp ik/y})/\mathbb{C}(y))=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}c\in\mathbb{C}^{\ast}\right\}.$
Here $\mathrm{Gal}(\mathbb{L}/\mathbb{K})$ denotes the differential Galois group, where $\mathbb{K}$ and $\mathbb{L}$ are differential fields such that $\mathbb{L}$ is a differential field extension of $\mathbb{K}$. The Stokes directions, on which $\mathrm{Re}\,(\mp ik/y)=0$, are given by $\arg y=0,\pi$. We have $\mathrm{Re}\,(-ik/y)<0$ and $\mathrm{Re}\,(ik/y)<0$ for $\arg y\in(0,\pi)$ and $(\pi,2\pi)$, respectively. So the singular directions $d_{\mp}$, which are bisectors of $(\tfrac{1}{2}\pi\mp\tfrac{1}{2}\pi,\tfrac{3}{2}\pi\mp\tfrac{1}{2}\pi)$, become $\arg y=\pi\mp\tfrac{1}{2}\pi$. See Fig. 2.
Let $a(k)\neq 0$. We take $\phi(1/y;k)$ and $\psi(1/y;-k)$ (resp. $\phi(1/y;-k)$ and $\psi(1/y;k)$) as solutions corresponding to the formal solution (4.3) with the upper (resp. lower) sign, in the sectors $\Sigma_{-}$ and $\Sigma_{+}$ containing $\mathrm{Re}\,y<0$ and $\mathrm{Re}\,y>0$, respectively, as well as the singular directions $d_{\pm}$. Near the singular direction $d_{-}$, we have
$\begin{pmatrix}\phi(1/y;k)&\phi(1/y;-k)\\ \phi^{\prime}(1/y;k)&\phi^{\prime}(1/y;-k)\end{pmatrix}=\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}$ (4.5)
by (1.7) and (2.3), so that the associated Stokes matrix is given by
$S_{-}=\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}.$
Near the singular direction $d_{+}$, we use (2.4) and (4.5) to obtain
$\begin{pmatrix}\psi(1/y;-k)&\psi(1/y;k)\\ \psi^{\prime}(1/y;-k)&\psi^{\prime}(1/y;k)\end{pmatrix}=\begin{pmatrix}\phi(1/y;-k)&\phi(1/y;k)\\ \phi^{\prime}(1/y;-k)&\phi^{\prime}(1/y;k)\end{pmatrix}\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix},$
so that the associated Stokes matrix is given by
$S_{+}=\begin{pmatrix}a(-k)&-b(-k)\\ -b(k)&a(k)\end{pmatrix}=S_{-}^{-1}.$
Ramis’ theorem [15, 22, 23] implies that the differential Galois group $G$ of (1.5) contains the Zariski closure of a group generated by the formal monodromy $M=\mathrm{id}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{\pm}$. In particular, $G$ is not finite, i.e., cases (i) and (ii) of Proposition 2.3 do not occur. Moreover, case (iv) of Proposition 2.3 does not occur since $a(k)\neq 0$. Hence, if Eq. (1.5) is integrable in the sense of differential Galois theory, then we have $b(k)$ or $b(-k)=0$ since either case (iii) or (v) of Proposition 2.3 occurs. So we have the following.
Proposition 4.1.
@@
Suppose that Eq. (1.5) is solvable by quadrature. Since by Proposition 2.1(iii) $a(k)=0$ only at discrete points, it follows from Proposition 4.1 that $b(k)$ or $b(-k)=0$ except at the discrete points. Hence, by the identity theorem (e.g., Theorem 3.2.6 of [2]), we have $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$ since by Proposition 2.1(i) $b(k)$ is analytic in a neighborhood of $\mathbb{R}^{\ast}$ in $\mathbb{C}^{\ast}$. Thus, we complete the proof. ∎
@@
In this section we give a proof of Theorem 1.6. The proof is mainly relied to the result of Kovacic [13], which is briefly described in Appendix A.
@@
Henceforth we assume that $m_{2}-m_{1}=1$. Obviously, case (c) of Proposition A.2 does not occur for $k\in\mathbb{R}^{\ast}$. So we only have to show that cases (a) and (b) of Proposition A.1 do not occur for some $k\in\mathbb{R}^{\ast}$. We rewrite (1.5) in the form (A.1) with
@@
We assume that case (a) of Proposition A.1 occurs as well as $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A2, we have $d_{c}\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ in (A.2) and
@@
since $\alpha=ik$ and $\beta=-\bar{u}_{1}$. We easily see that $d_{c}\not\in\mathbb{N}_{0}$ when $k$ changes slightly even if $d_{c}\in\mathbb{N}_{0}$for some $k\in\mathbb{R}^{\ast}$ This implies that case (a) of Proposition A.1 does not occur in an open interval of $k\in\mathbb{R}^{\ast}$ since $\mathscr{S}$ is finite so that the number of a finite sequence $c$ of the symbols $\pm$ with $d_{c}\in\mathbb{N}_{0}$ is finite.
We next assume that case (b) of Proposition A.1 occurs for an open interval of $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A3, we have $d_{e}\in\mathbb{N}_{0}$ in (A.3) and there exists a monic polynomial $P(x)$ of order $d_{e}$ satisfying (A.5). Since
@@
we easily see that $d_{e}\neq 0$. Let $d=d_{e}\geq 1$. Since Eq. (A.5) holds, we see that
@@
since the functions given by (5.2) are linearly dependent on the interval of $k$. Substituting (5.3) and (5.4) into (A.5), we obtain
@@
and let $r(x)=-k^{2}-u(x)$. Equation (1.5) satisfies $m_{2}-m_{1}=1$ and only has singularities of order two at $x=0$ and $1$. Following the recipe of Appendix A.2, we have $\beta=5/16$ and
@@
Hence, case (a) of Proposition A.1 does not occur for any $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix A.3, we have $\beta=5/16$ and
@@
Substituting $P(x)=x+p_{0}$ into (A.5), we obtain
@@
Hence, case (b) of Proposition A.1 holds with $p_{0}=-\tfrac{1}{4}$, so that Eq. (1.5) is solvable by quadrature, if and only if $k=\pm\tfrac{2}{3}\sqrt{3}$.
@@
so that the integrand in (A.6) satisfies
@@ Acknowledgements
Appendix A Kovacic’s Result

### v2

@@ 1. Introduction
  • (ii)
    Statements similar to those of Theorems 1.3 and 1.4 are proven for the ZS systems in [30].
@@
The outline of this paper is as follows. In Section 2 we collect preliminary results for proving Theorems 1.3 and 1.4. We give proofs of Theorems 1.3, 1.4 and 1.6 in Sections 3, 4 and 5, respectively. Fundamental information on local differential Galois groups near irregular singularities, which is needed in Section 4, is provided in Appendix A, and the result of Kovacic [13], which plays a crucial role in Section 5, is briefly reviewed in Appendix B.
@@ 2. Preliminary Results
The conclusion in Proposition 2.1(iv) holds under a weaker condition. For example, if $\lim_{k\to 0}b(k)b(-k)=b_{0}$ $(\neq-1)$, then it holds.
@@
The system (1.5) is solvable by quadrature unless $G$ is of type (vi). Proposition 2.3 plays a key role in the proof of Theorem 1.4 in Section 4.
@@
where $\tilde{u}(y)=u(1/y)$. Since $(k^{2}-\tilde{u}(y))/y^{4}$ has a pole of fourth order at $y=0$, we see by a standard result on higher-order scalar linear differential equations (see, e.g., Theorem 7 of [6]) that Eq. (1.5) has an irregular singularity at infinity $x=\infty$. To obtain a necessary condition for its integrabilty, we compute its formal monodromy, exponential torus and Stokes matrices around $y=0$ for (4.1). See Appendix A for necessary information, and see also [21, 23, 28] for more details.
@@
$\frac{\mathrm{d}\eta}{\mathrm{d}y}+\eta^{2}+\frac{2}{y}\eta+\frac{k^{2}-\tilde{u}(y)}{y^{4}}=0.$ (4.2)
@@
Substitution of $ye^{2\pi i}$ into (4.4) yields
$V(ye^{i\theta})=\tilde{V}(ye^{2\pi i})e^{Q(ye^{2\pi i})},$
@@
$\hat{M}=V(y)^{-1}V(ye^{2\pi i})=\mathrm{id}$
by $V(ye^{2\pi i})=V(y)\hat{M}$. The exponential torus is given by
$\mathscr{T}=\mathscr{T}:=\mathrm{Gal}\left(\mathbb{C}((y))(e^{ik/y},e^{-ik/y})/\mathbb{C}((y))\right)=\left\{\begin{pmatrix}c&0\\ 0&c^{-1}\end{pmatrix}\bigg{|}c\in\mathbb{C}^{\ast}\right\}.$
Here $\mathrm{Gal}(\mathbb{L}/\mathbb{K})$ denotes the differential Galois group, where $\mathbb{K}$ and $\mathbb{L}$ are differential fields such that $\mathbb{L}$ is a differential field extension of $\mathbb{K}$. The Stokes directions, on which $\mathrm{Re}\,(\mp ik/y)=0$, are given by $\arg y=0,\pi$. We have $\mathrm{Re}\,(-ik/y)<0$ and $\mathrm{Re}\,(ik/y)<0$ for $\arg y\in(0,\pi)$ and $(\pi,2\pi)$, respectively. So the singular directions $d_{\mp}$, which are bisectors of $(\tfrac{1}{2}\pi\mp\tfrac{1}{2}\pi,\tfrac{3}{2}\pi\mp\tfrac{1}{2}\pi)$, become $\arg y=\pi\mp\tfrac{1}{2}\pi$. See Fig. 2. By Proposition  A.1, we can write the Stokes matrices as
$S_{-}=\begin{pmatrix}1&\alpha_{-}\\ 0&1\end{pmatrix}\quad\mbox{and}\quad S_{+}=\begin{pmatrix}1&0\\ \alpha_{+}&1\end{pmatrix}$ (4.5)
for the singular directions $d_{-}$ and $d_{+}$, respectively, where $\alpha_{\pm}$ are constants which may be zero.
Suppose that $a(k),a(-k)\neq 0$. Let
$\Phi(y;k)=\begin{pmatrix}\phi(y^{-1};k)&\phi(y^{-1};-k)\\ -y^{-2}\phi^{\prime}(y^{-1};k)&-y^{-2}\phi^{\prime}(y^{-1};-k)\end{pmatrix},$
$\Psi(y;k)=\begin{pmatrix}\psi(y^{-1};-k)&\psi(y^{-1};k)\\ -y^{-2}\psi^{\prime}(y^{-1};-k)&-y^{-2}\psi^{\prime}(y^{-1};k)\end{pmatrix},$
which are fundamental matrices to (4.1). By (1.7) and (2.3),
$\Phi(y;k)=\Psi(y;k)B(k),\quad B(k)=\begin{pmatrix}a(k)&b(-k)\\ b(k)&a(-k)\end{pmatrix}.$ (4.6)
Lemma 4.1.
If $\alpha_{-}$ and $\alpha_{+}=0$, then $b(-k)$ and $b(k)=0$, respectively.
Proof.
We prove that if $\alpha_{-}=0$, then $b(-k)=0$. The remaining part can be proven similarly.
First, by (1.4), we can write the formal fundamental matrix as
$V(y)=\left(\begin{array}[]{cc}\begin{matrix}*\\ *\end{matrix}\quad\Phi(y;k)\begin{pmatrix}0\\ c_{-}\end{pmatrix}\end{array}\right)\quad\mbox{and}\quad\left(\begin{array}[]{cc}\begin{matrix}*\\ *\end{matrix}\quad\Psi(y;k)\begin{pmatrix}0\\ c_{+}\end{pmatrix}\end{array}\right)$
in the directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ near $d_{-}$ in $\Sigma_{-}$ and $\Sigma_{+}$, respectively, where $c_{\pm}\in\mathbb{C}$ are nonzero constants. If $\alpha_{-}=0$, then we use (4.5) and (4.6) to obtain
$b(-k)c_{-}=0,\quad a(-k)c_{-}-c_{+}=0,$
by the definition of the Stokes matrix. Thus, we obtain the desired result. ∎
Theorem A.2 implies that the differential Galois group $G$ of (1.5) contains the Zariski closure of a group generated by the formal monodromy $\hat{M}=\mathrm{id}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{\pm}$. In particular, $G$ is not finite, i.e., cases (i) and (ii) of Proposition 2.3 do not occur. Moreover, case (iv) of Proposition 2.3 does not occur since $a(k)\neq 0$. Hence, if Eq. (1.5) is integrable in the sense of differential Galois theory, then we have have $\alpha_{-}$ or $\alpha_{+}=0$, so that by Lemma 4.1 $b(k)$ or $b(-k)=0$, since either case (iii) or (v) of Proposition 2.3 occurs. So we have the following.
Lemma 4.2.
@@
Suppose that Eq. (1.5) is solvable by quadrature. Since by Proposition 2.1(iii) $a(k)=0$ only at discrete points, it follows from Lemma 4.2 that $b(k)$ or $b(-k)=0$ except at the discrete points. Hence, by the identity theorem (e.g., Theorem 3.2.6 of [2]), we have $b(k)=0$ for any $k\in\mathbb{R}^{\ast}$ since by Proposition 2.1(i) $b(k)$ is analytic in a neighborhood of $\mathbb{R}^{\ast}$ in $\mathbb{C}^{\ast}$. Thus, we complete the proof. ∎
@@
In this section we give a proof of Theorem 1.6. The proof is mainly relied to the result of Kovacic [13], which is briefly described in Appendix B.
@@
Henceforth we assume that $m_{2}-m_{1}=1$. Obviously, case (c) of Proposition B.2 does not occur for $k\in\mathbb{R}^{\ast}$. So we only have to show that cases (a) and (b) of Proposition B.1 do not occur for some $k\in\mathbb{R}^{\ast}$. We rewrite (1.5) in the form (B.1) with
@@
We assume that case (a) of Proposition B.1 occurs as well as $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix B2, we have $d_{c}\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ in (B.2) and
@@
since $\alpha=ik$ and $\beta=-\bar{u}_{1}$. We easily see that $d_{c}\not\in\mathbb{N}_{0}$ when $k$ changes slightly even if $d_{c}\in\mathbb{N}_{0}$for some $k\in\mathbb{R}^{\ast}$ This implies that case (a) of Proposition B.1 does not occur in an open interval of $k\in\mathbb{R}^{\ast}$ since $\mathscr{S}$ is finite so that the number of a finite sequence $c$ of the symbols $\pm$ with $d_{c}\in\mathbb{N}_{0}$ is finite.
We next assume that case (b) of Proposition B.1 occurs for an open interval of $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix B3, we have $d_{e}\in\mathbb{N}_{0}$ in (B.3) and there exists a monic polynomial $P(x)$ of order $d_{e}$ satisfying (B.5). Since
@@
we easily see that $d_{e}\neq 0$. Let $d=d_{e}\geq 1$. Since Eq. (B.5) holds, we see that
@@
since the functions given by (5.2) are linearly dependent on the interval of $k$. Substituting (5.3) and (5.4) into (B.5), we obtain
@@
and let $r(x)=-k^{2}-u(x)$. Equation (1.5) satisfies $m_{2}-m_{1}=1$ and only has singularities of order two at $x=0$ and $1$. Following the recipe of Appendix B.2, we have $\beta=5/16$ and
@@
Hence, case (a) of Proposition B.1 does not occur for any $k\in\mathbb{R}^{\ast}$. Following the recipe of Appendix B.3, we have $\beta=5/16$ and
@@
Substituting $P(x)=x+p_{0}$ into (B.5), we obtain
@@
Hence, case (b) of Proposition B.1 holds with $p_{0}=-\tfrac{1}{4}$, so that Eq. (1.5) is solvable by quadrature, if and only if $k=\pm\tfrac{2}{3}\sqrt{3}$.
@@
so that the integrand in (B.6) satisfies
@@ Acknowledgements
Appendix A Local Differential Galois Groups near Irregular Singularities
In this appendix we review basic information on local differential Galois groups, which consist of formal monodromies, exponential tori and Stokes matrices, near irregularity singular points. See [21, 23, 28] for the details. Recall some terminologies of the differential Galois theory in Section 3 of [29] (see also [10, 23]) if necessary in the following.
Consider $n$-dimensional linear systems of the form
$\frac{\mathrm{d}\zeta}{\mathrm{d}y}=A(y)\zeta$ (A.1)
with an irregular singularity at $y=0$ in $\mathbb{C}$, where $A(y)\in\mathrm{gl}(n,\mathbb{C}((y)))$, and $\mathbb{C}((y))$ denotes the field of the formal Laurent series over $\mathbb{C}$. In a standard way, Eq. (4.1) can be written in the form (A.1) easily. The linear system (A.1) has a formal fundamental matrix of the form
$Y(y)=\hat{Y}(y)y^{L}e^{Q(1/z)}$ (A.2)
like (4.4), where
  • $y^{\ell}=z$ for some $\ell=1,\ldots,n$;
  • $\hat{Y}(y)\in\mathrm{GL}(n;\mathbb{C}((y)))$;
  • $L\in\mathrm{gl}(n;\mathbb{C})$;
  • $Q(1/z)=\operatorname{diag}(q_{1}(z),\ldots,q_{n}(z))$, $q_{j}(z)\in\frac{1}{z}\mathbb{C}[\frac{1}{z}]$
(see, e.g., [16]). So we obtain a formal Picard-Vessiot extension $\mathbb{L}\subset\mathbb{K}=\mathbb{C}((y))$, where $\mathbb{L}$ is a differential field generated by entries of $Y(y)$ and having $\mathbb{C}$ as its field of constants.
Let $\hat{\mu}$ be a $\mathbb{K}$-automorphism of $\mathbb{L}$ satisfying
  • $\hat{\mu}(y^{\alpha_{j}})=e^{2\pi i\alpha_{j}}y^{\alpha_{j}}$ for any $j=1,\ldots,n$;
  • $\hat{\mu}(\log y)=\log y+2\pi i$;
  • $\hat{\mu}(\exp(q_{j}(y^{1/\ell})))=\exp(q_{j}(e^{2\pi i/\ell}y^{1/\ell})$,
where $\alpha_{j}$, $j=1,\ldots,n$, denote the eigenvalues of $L$. Since $\hat{\mu}$ commutes the differentiation, $\hat{\mu}(Y(y))$ is also a formal fundamental matrix. Hence, there exists a matrix $\hat{M}\in\mathrm{GL}(n;\mathbb{C})$ such that
$\hat{\mu}(Y(y))=Y(y)\hat{M}.$
We call $\hat{M}$ the formal monodromy of (A.1). Thus, $\hat{M}$ can be computed by substituting $ye^{i\theta}$ into $Y(y)$, as in Section 4.
Let $\mathbb{E}$ be a differential field extension of $\mathbb{K}$ such that $\mathbb{L}=\mathbb{E}(e^{q_{1}(y^{1/\ell})},\ldots,e^{q_{n}(y^{1/\ell})})$. Then $\mathscr{T}=\mathrm{Gal}(\mathbb{L}/\mathbb{E})$ is called the exponential torus of (A.1). Thus, $\mathscr{T}$ is the group of all $\mathbb{E}$-automorphisms of $\mathbb{L}$.
We next state the definition of the Stokes matrix for (4.2). Some preparations are needed to this end. Let $q(z)\in z^{-1}\mathbb{C}[z^{-1}]$, so that $q(z^{-1})$ is a polynomial of $z$. Let $k$ be the degree of $q(z^{-1})$ and let $q(z)=c_{0}z^{-k}+\ldots$, where $c_{0}\in\mathbb{C}$ is a constant. The direction $d$ determined by $\mathrm{Re}\,(c_{0}y^{-k/\ell})=0$, in which $e^{q(z)}$ changes its behavior, is called the Stokes direction for $q(z)$. Let $d_{1},d_{2}$ be consecutive Stokes directions. We say that $(d_{1},d_{2})$ is a negative Stokes pair if $\mathrm{Re}\,(c_{0}y^{-k/\ell})<0$ for $\arg(y)\in(d_{1},d_{2})$. We refer to the bisector a negative Stokes pair as a singular direction of $q(z)$.
Figure 3. Singular direction $d$ and the two directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$.
We now consider the formal fundamental matrix (A.2). Let $d$ be a singular direction of $q_{j}(z)-q_{l}(z)$ for some $j,l=1,\ldots,n$ with $j\neq l$. Choose two directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ counterclockwise nearby such that the associated sectors overlap and contain $d$. See Fig. 3. Let $\hat{Y}_{\mathrm{L}}(y)$ and $\hat{Y}_{\mathrm{R}}(y)$ be the multisums of $\hat{Y}(y)$ in the directions $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$, respectively. Then there exists $S_{jl}\in\mathrm{GL}(n;\mathbb{C})$ such that
$\hat{Y}_{\mathrm{L}}(y)y^{L}e^{Q(1/z)}=\hat{Y}_{\mathrm{R}}(y)y^{L}e^{Q(1/z)}S_{jl}$ (A.3)
since Eq. (A.2) gives a formal fundamental matrix for both cases. The matrix $S_{jl}$ is independent of the choice of $d_{\mathrm{L}}$ and $d_{\mathrm{R}}$ and called a Stokes matrix for the singular direction $d$. We have the following on the Stokes matrices (see, e.g., Proposition 2.37 of [21] for its proofs).
Proposition A.1.
The Stokes matrices are unipotent such that all the diagonal elements are one and the $(j^{\prime},l^{\prime})$-element is zero if $d$ is a singular direction for $q_{j^{\prime}}(z)-q_{l^{\prime}}(z)$.
Finally, we state one of the most important results on local differential Galois groups near irregular singularities as follows.
Theorem A.2 (Ramis [24, 25]).
The local differential Galois group of (4.2) around the irregular singular point $y=0$ is the Zariski closure of the group generated the formal monodromy $\hat{M}$, the exponential torus $\mathscr{T}$ and the Stokes matrices $S_{jl}$, $j,l=1,\ldots,n$ with $j\neq l$.
See also [15, 23] for a proof of Theorem A.2. Theorem A.2 is regarded as a generalization of Schlesinger’s theorem (see, e.g., [23, 29]) for regular singularities to irregular ones.
Appendix B Kovacic’s Result